OpenStudy (anonymous):

Solve -7 < x + 2 < 4. Answer -9 < x < 2 -5 < x < 6 x < 2 or x > -9 no solution

OpenStudy (anonymous):

-9 < x < 2, because -9+2=7 and 2+2=4

OpenStudy (anonymous):

you have to us BEDMAS - brackets, exponents, division, multiplication, additition, subtraction

OpenStudy (johnweldon1993):

-7 < x + 2 < 4. You just want to isolate 'x' because that is what we are solving for....so to isolate 'x' here....we need to subtract 2 from everything -7 < x + 2 < 4. -2 -2 -2 *see how +2 -2 = 0 so it cancels out* so we have -9 < x < 2

OpenStudy (anonymous):

okay thanks

OpenStudy (johnweldon1993):

No problem!